without using finance calculator,
Refer to Johan's investment account, i.e. the monthly deposits made into his investment account:

Year Monthly deposits
2017 SR 500
2018 SR 400
2019 SR 700

Find the value of his investment at the end of 2019 if the investment rate was 12% compounded monthly. Find also the interest earned.
1
Expert's answer
2020-05-28T10:59:51-0400

1)"S=500*\frac{1.12^{12}-1}{1.12-1}+400*\frac{1.12^{12}-1}{1.12-1}+700*\frac{1.12^{12}-1}{1.12-1}=38,612.8"

2)"S^*=38,612.8-500*12-400*12-700*12=19,412.8"